An object that is projected straight downward
with initial velocity v feet per second travels a
distance s=vt+16t^2, where t = time in seconds.
If Ramón is standing on a balcony 84 feet above
the ground and throws a penny straight down
with an initial velocity of 10 feet per second, in
how many seconds will it reach the ground?

Answer by ankor@dixie-net.com(22740)

An object that is projected straight downward with initial velocity v feet per
second travels a distance s=vt+16t^2, where t = time in seconds.
If Ramón is standing on a balcony 84 feet above
the ground and throws a penny straight down
with an initial velocity of 10 feet per second, in
how many seconds will it reach the ground?
:
The equation for this would be:
(negative means down)
-16t^2 - 10t + 84 = 0
solve this using the quadratic equation, the positive solution: t = 2 sec